Cynthia is planning a party. For entertainment, she has designed a game that involves spinning two spinners. If the sum of the numbers on the spinners is 10 or greater, the guests can choose a prize from the basket of candy bars. If the sum is less than 1o, then the guests will be thrown in the pool. Determine the probability of getting tossed in the pool.
Spinner A: 4 parts, 3/4 is 1, 1/4 is 3.
Spinner B: 3 parts, 1/3 is 5, 1/3 is 7, 1/3 is 9.

Answer by fractalier(6550)

The combinations of the two spinners are
1 and 5
1 and 7
1 and 9
3 and 5
3 and 7
3 and 9
The first, second and fourth cases get you thrown in the pool.
Thus let's add up those probabilities...
1 and 5 ==> 3/4*1/3 = 1/4
1 and 7 ==> 3/4*1/3 = 1/4
3 and 5 ==> 1/4*1/3 = 1/12
Add 'em and get 7/12, the chances you're going swimming...